# Automated Defects Detection and Fix in Logging Statement

### v1

@@ 2.3.3. Temporal inconsistency
The tense of the logging text is inconsistent with the actual situation. The tense used in logging statements should align with the temporal context of the events. This adherence ensures clarity for tracking system events. In Fig. 1(c), the original logging statement uses the present continuous tense “Creating”, which implies that the action of creating the exec probe is still ongoing. However, the logging statement is placed immediately before the return statement, indicating that the creation of the probe has been completed by the time of the logging statement execution.
Figure 2. The overview framework of LogFixer.
@@ 2.3.4. Readability issues
The logging statement with spelling errors or abusive capitalization can undermine their readability. Eliminating these errors is crucial to prevent misunderstandings and ensuring that logs serve as clear and effective communication tools for developers and operators. Taking Fig. 1(d) as an example, ”Non-optimial” is a typo error.
Summary: The study uncovers four categories of factual defects (accounting for 51.6% LCC) that can affect the quality of logging statements.
@@ 3.1. Overview
To this end, we propose LogFixer, a framework for automatically detecting and updating defective logging statements. Fig. 2 illustrates the high-level view of LogFixer, which operates in two stages: offline fine-tuning and online updating. During the offline stage, the goal is to develop a defective logging statement detector. We first synthesize the defective logging statements with corpus collected from well-maintained repositories, ensuring high-quality code and maintenance standards. Then we use the synthetic data to fine-tune a similarity-based logging defect detector. The online phase involves two steps: detecting defective logging statements and updating them. We first use the fine-tuned detector to determine the $type_{s}$. Based on the defect information, we apply the analyzed knowledge into the LLM-based updater for logging statement update.
@@ 3.2.3. LLM-based mutation
For defects requiring semantic mutation (statement-code inconsistency and static-dynamic inconsistency), using rule-based or heuristic methods can be challenging. Therefore, we employ LLM-based techniques to synthesize defective logging statements. These LLMs, trained on vast amounts of code, have been shown to handle code-related downstream tasks effectively (Zhang et al., 2024a; Xia et al., 2023; Yang et al., 2023). We use GPT-4o (OpenAI, 2024) to mutate the semantics for static content or dynamic variables. Specifically, for statement-code inconsistency, we prompt the LLM to carefully analyze the surrounding code and then alter the static content of logging statements to create semantic inconsistencies (e.g., recording opened while the actual event is close). For static-dynamic inconsistency, we instruct the LLM to randomly mutate the static content or dynamic variables to generate potential mismatches (e.g., recording exc while a directory path dir is expected).
@@ 3.3. Logging Statement Defects Detection
Considering that only a small portion of logging statements require update, and scanning all methods in a repository with LLM is resource-intensive, LogFixer employs a similarity-based classifier as the detector. This detector identifies defective logging statements and determines the types of defects. The introduction of the fine-tuned detector significantly reduces costs and enables subsequent LLM components to update the logging statements in a more efficient and targeted manner.
Specifically, this phase is responsible for handling a multi-class classification task. Given a logging statement $s_{original}$ with its associated context (i.e., code snippet), the goal is to determine whether $s_{original}$ has a specific type of defect. We resolve the problem through the following steps:
@@ 3.3.2. Similarity-based multiple classification
3.4. Defective Logging Statement Update
Figure 3. The recommendation framework on a static-dynamic inconsistency case.
Following the defect classification phase, where potential defective logging statements are identified, we proceed with a multi-expert collaboration framework to repair these statements, inspired by the cognitive synergy (Goertzel, 2009). In particular, we employ two LLM-based experts: a checker and an updater. Fig. 3 depicts a synergy between two experts, as well as the specific duty, and the prompt for each expert. The checker’s firstly validates the classification results from the detector, and then employs LLMs to provide supplementary semantic information. Afterward, the updater is responsible for implementing logging statement updates according to the advice provided by the checker.
@@ 3.4.1. Checker
The role of checker is to reduce the false positives brought by the similarity-based detector, where non-defective logging statements could be mistakenly marked as requiring update due to the extremely imbalanced data in real-world situations. To address this issue, we have designed a checker with two objectives. The first objective is to determine whether the detector’s decisions on updating logging statements are necessary, aiming to eliminate false positives. The second objective is to leverage the extensive knowledge of the LLM to offer additional semantic insights, thereby enhancing the effectiveness of subsequent experts.
@@ 3.4.2. Updater
We utilize log-centric changes as supplementary knowledge to guide the LLM in learning the past update patterns from maintainers. We apply a defect type-aware strategy to select proper historical LCC examples for the updater’s reference. In specific, for defect types sensitive to project characteristics (e.g., temporal relation inconsistency and readability issues), we exclusively select the LCCs from the same project to maintain the consistency in the project’s logging style (Li et al., 2024a). In contrast, for defects related to semantic information inconsistency that are unaffected by logging style and inter-project variations, we incorporate LCCs from multiple projects for the updater. We employ the BM25 (Robertson et al., 2009) ranking algorithm to identify LCCs that are the most similar to the target logging statements.
@@ 4.1. Fine-tuning Dataset
To construct high-quality data for fine-tuning the detector, we apply the technique mentioned in 3.2 to generate 5K defective instances for each type. Meanwhile, we consider logging statements from well-maintained repositories as defect-free because they adhere to stringent community guidelines and thorough maintenance protocols. Specifically, we select well-maintained repositories from the target repositories based on the following criteria:
@@ 4.2. Implementation Details
During LLM-based defective logging statement synthesis, we use the GPT-4o (OpenAI, 2024) API provided by Openai as the default LLM model. We use UniXcoder-base as the tokenization and embedding tool. We use Adam (Kingma and Ba, 2017) as the optimizer and set the learning rate to 5e-5, Adam epsilon to 1e-8, and dropout rate to 0.1 when fine-tuning the similarity-based classification model. We fine-tune the model for 10 epochs and set the max token number to 1024. We run them on a server with Dual Intel(R) Gold 6326 CPU, 512GB physical memory, and 4x NVIDIA RTX 3090 GPU. The OS version is Ubuntu 18.04.
@@ 5. Evaluation
We evaluate LogFixer by answering the following research questions:
RQ1: How effective is LogFixer at detecting and enhancing defective logging statements?
RQ2: How do different components affect the performance of LogFixer?
RQ3: How generalizable is LogFixer for different LLM backbones?
RQ4: How is the practitioners’ feedback on LogFixer?
@@ 5.1. Evaluation Dataset
To evaluate the performance of LogFixer, we conducted experiments on three distinct datasets.
@@ 5.1.1. Existing logpatches
We utilize the data annotated in the manual study (2.2) as existing logpatches. Each sample forms a pair, consisting of an original defective logging statement and its corrected version, culminating in a total of 258 defective pairs. Additionally, we sample 258 non-defect logging statements from well-maintained repositories, which brings our total case count to 516.. Note that we exclude these data from the training set.
@@ 5.1.3. New repositories
To evaluate the effectiveness of LogFixer in addressing previously unidentified defects, we randomly sampled 100K log-code pairs outside the collected well-maintained dataset for human evaluation. These samples were deliberately excluded from all previous processes, thereby guaranteeing that the evaluation was conducted on entirely fresh data.
@@ 5.2. Metrics
For the detection ability of LogFixer, we use Precision, Recall, and F1-macro to measure multiple classification performance. Formally, $Precision=\frac{TP}{TP+FP}$, $Recall=\frac{TP}{TP+FN}$, $F_{1}=2*(\frac{Precision*Recall}{Precision+Recall})$, and $F_{1}^{macro}=\frac{1}{N}\sum_{i=1}^{N}F_{1}^{i}$. We use $F_{1}^{macro}$ to give equal weight to each class, avoiding the issue of inflating the weight of the class with more instances.
For the updating ability, we use the same metrics as logging statement generation task. Different metrics are used to evaluate different parts of the logging statements.
(1) Static content. In alignment with recent research efforts (Li et al., 2023b; Mastropaolo et al., 2022), the quality of updated logging texts is evaluated using two established metrics from the field of machine translation: BLEU-K (where K = 1, 2, 4) and ROUGE-K (where K = 1, 2, L), to measure the overlap in terms of K-grams between the updated and actual logging texts.
(2) Dynamic variables. Following the recent work (Li et al., 2023b), we utilize special Precision, Recall, and F1 to assess the accuracy of updated logging variables For each logging statement, let $S_{ud}$ represent the variables in the LogFixer updates, and $S_{gt}$ denote the variables in the actual logging statement. We calculate and report the proportion of correctly updated variables ($precision$=$\frac{S_{ud}\cap S_{gt}}{S_{ud}}$), the proportion of actual variables that are correctly predicted by the model ($recall$=$\frac{S_{ud}\cap S_{gt}}{S_{gt}}$), and their harmonic mean ($F_{1}$=$2*\frac{Precision*Recall}{Precision+Recall}$).
@@
5.3. Baseline
Table 3. The detection results on synthetic data and existing logpatches.
@@ 5.3. Baseline
CodeT5+ 0.314 0.062 0.096
LogFixer 0.725 0.552 0.625
@@
CodeT5+ 0.366 0.075 0.115
LogFixer 0.925 0.673 0.776
To provide a comprehensive evaluation, we compare the detection ability and update ability of LogFixer with representative methods, with details as below.
For the detection ability, we pre-trained a CodeT5+ (Wang et al., 2023) model based on the approach proposed by CMI-Finder (Bouzenia and Pradel, 2023) to serve as a baseline for comparison, which has shown good performance in the binary classification task of message-condition inconsistency.
@@
5.4. RQ1: Effectiveness of LogFixer
Table 4. The updating result on existing logpatches and synthetic data.
{NiceTabular}
l—llllll—lll Static texts Dynamic variables
BLEU-1 BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L Precision Recall F1
Existing logpatches
GPT4o-Instruct 0.690 (4.90$\uparrow$) 0.587 (12.86$\uparrow$) 0.268 (0.94$\uparrow$) 0.746 (13.01$\uparrow$) 0.593 (17.44$\uparrow$) 0.745 (12.96$\uparrow$) 0.272 (0.13$\uparrow$) 0.256 (1.59$\uparrow$) 0.264 (1.15$\uparrow$)
GPT4o-ICL 0.681 (2.14$\uparrow$) 0.588 (13.08$\uparrow$) 0.262 (0.14$\uparrow$) 0.750 (14.38$\uparrow$) 0.603 (19.47$\uparrow$) 0.748 (13.99$\uparrow$) 0.300 (3.97$\uparrow$) 0.307 (21.82$\uparrow$) 0.304 (16.47$\uparrow$)
GPT4o-DET 0.701 (8.28$\uparrow$) 0.620 (19.83$\uparrow$) 0.327 (8.93$\uparrow$) 0.781 (25.00$\uparrow$) 0.650 (29.01$\uparrow$) 0.779 (24.57$\uparrow$) 0.315 (6.03$\uparrow$) 0.296 (17.46$\uparrow$) 0.305 (16.85$\uparrow$)
LogFixer 0.799 (38.34$\uparrow$) 0.721 (41.13$\uparrow$) 0.384 (16.64$\uparrow$) 0.849 (48.28$\uparrow$) 0.738 (46.85$\uparrow$) 0.848 (48.12$\uparrow$) 0.439 (23.04$\uparrow$) 0.451 (26.60$\uparrow$) 0.445 (24.90$\uparrow$)
Synthetic data
GPT4o-Instruct 0.546 (-7.32$\downarrow$) 0.465 (-7.86$\downarrow$) 0.316 (-2.54$\downarrow$) 0.634 (-6.08$\downarrow$) 0.499 (-9.15$\downarrow$) 0.632 (-6.35$\downarrow$) 0.696 (14.85$\uparrow$) 0.707 (42.20$\uparrow$) 0.702 (32.57$\uparrow$)
GPT4o-ICL 0.545 (-7.56$\downarrow$) 0.468 (-7.25$\downarrow$) 0.313 (-3.00$\downarrow$) 0.639 (-4.64$\downarrow$) 0.508 (-7.19$\downarrow$) 0.637 (-4.91$\downarrow$) 0.683 (11.20$\uparrow$) 0.664 (33.73$\uparrow$) 0.673 (26.02$\uparrow$)
GPT4o-DET 0.640 (14.89$\uparrow$) 0.579 (15.12$\uparrow$) 0.371 (5.69$\uparrow$) 0.715 (17.39$\uparrow$) 0.636 (20.69$\uparrow$) 0.714 (17.34$\uparrow$) 0.752 (30.53$\uparrow$) 0.761 (52.85$\uparrow$) 0.756 (44.79$\uparrow$)
LogFixer 0.659 (19.39$\uparrow$) 0.592 (17.74$\uparrow$) 0.436 (15.44$\uparrow$) 0.753 (28.40$\uparrow$) 0.648 (23.31$\uparrow$) 0.751 (28.03$\uparrow$) 0.795 (42.57$\uparrow$) 0.778 (56.21$\uparrow$) 0.787 (51.80$\uparrow$)
Approach. The effectiveness of LogFixer should be evaluated from two aspects: first, the detection capability, specifically the capability to identify defective logging statements; second, the updating capability, which pertains to the similarity between the revised logging statements (i.e., prediction) and the actual ground truth (i.e., repository update history).
Result. Table 3 shows the detection results with different classification models. Overall, our proposed detector significantly outperforms CodeT5+ across all metrics on both existing logpatches and synthetic datasets. This could be because the T5 architecture might struggle with complex code comprehension, as it lacks the utilization of structural features in the programming language (e.g., AST).
@@ 5.4. RQ1: Effectiveness of LogFixer
Table 5.4 illustrates the updating performance of static content. Compared to the three baselines, LogFixer demonstrates significant performance improvements across various metrics on two datasets. Specifically, for GPT4o-Instruct and GPT4o-ICL, which lack defect information, the improvement on the logpatches dataset is minimal (0.14$\uparrow$ IC on BLEU-4), and there is even a regression (3.00$\downarrow$) on synthetic data. However, GPT4o-DET, which includes defect information, shows a certain level of improvement, highlighting the importance of defect information.

### v2

@@ 2.3.3. Temporal inconsistency
To address temporal inconsistency, we adhere to the definition proposed by Ding et al. (Ding et al., 2023). This definition characterizes temporal inconsistency as arising when there is a mismatch between the logical temporal relations, which pertain to the sequence in which a logging statement is executed in relation to its associated source code (i.e., what the actual order is in the code), and the semantic temporal relations, which can be deduced from the semantic interpretation of the logging text (i.e., what the order is inferred from the generated logs). The tense used in logging statements should align with the temporal context of the events. This adherence ensures clarity for tracking system events. In Fig. 1(c), the original logging statement uses the present continuous tense “Creating”, which implies that the action of creating the exec probe is still ongoing. However, the logging statement is placed immediately before the return statement, indicating that the creation of the probe has been completed by the time of the logging statement execution.
@@ 2.3.4. Readability issues
Readability issues are linguistic defects that impairing logging message clarity, which includes spelling errors and capitalization errors. Eliminating these errors is crucial to prevent misunderstandings and ensure that logs serve as clear and effective communication tools for developers and operators. Taking Fig. 1(d) as an example, ”Non-optimial” is a typo error. The readability issue of logging statements is significant since many logs generated by logging statements are processed by automated tools for downstream tasks like anomaly detection and root cause analysis. While current models used in these tools can often tolerate minor typos, such defects can still degrade the performance of embeddings derived from logs, ultimately impacting the accuracy of automated analyses.
Summary: The study uncovers four categories of factual defects (statement-code inconsistency, static-dynamic inconsistency, temporal inconsistency, and readability issues), which account for 51.6% log-centric commits.
@@ 3.1. Overview
To this end, we propose LogUpdater, a framework for automatically detecting and updating defective logging statements. Fig. 2 illustrates the high-level view of LogUpdater, which operates in two stages: offline fine-tuning and online updating. During the offline stage, the goal is to develop a defective logging statement detector. We first synthesize the defective logging statements with corpus collected from well-maintained repositories, ensuring high-quality code and maintenance standards. Then we use the synthetic data to fine-tune a similarity-based logging defect detector. The online phase involves two steps: detecting defective logging statements and updating them. We first use the fine-tuned detector to determine the $type_{s}$. Based on the defect information, we apply the analyzed knowledge into the LLM-based updater for logging statement update.
Figure 2. The overview framework of LogUpdater.
@@ 3.2.1. Mutation of words
Listing 2: A word mutation case for readability issue.
@@ 3.2.2. Mutation of tense
Listing 3: A tense mutation case for temporal inconsistency.
@@ 3.2.3. LLM-based mutation
For defects requiring semantic mutation (statement-code inconsistency and static-dynamic inconsistency), using rule-based or heuristic methods can be challenging. Therefore, we employ LLM-based techniques to synthesize defective logging statements. These LLMs, trained on vast amounts of code, have been shown to handle code-related downstream tasks effectively (Zhang et al., 2024a; Xia et al., 2023; Yang et al., 2023). We use GPT-4o (OpenAI, 2024) to mutate the semantics for static content or dynamic variables. Specifically, for statement-code inconsistency, we prompt the LLM to carefully analyze the surrounding code and then alter the static content of logging statements to create semantic inconsistencies (e.g., recording opened while the actual event is close). For static-dynamic inconsistency, we instruct the LLM to randomly mutate the static content or dynamic variables to generate potential mismatches (e.g., recording exc while a directory path dir is expected). The prompt details are shown in (Logupdater, 2025b).
Listing 4: A LLM-based mutation case for statement-code inconsistency.
@@
Listing 5: A LLM-based mutation case for static-dynamic inconsistency.
@@ 3.3. Logging Statement Defects Detection
Considering that only a small portion of logging statements require update, and scanning all methods in a repository with LLM is resource-intensive, LogUpdater employs a similarity-based classifier as the detector. This detector identifies defective logging statements and determines the types of defects. The introduction of the fine-tuned detector significantly reduces costs and enables subsequent LLM components to update the logging statements in a more efficient and targeted manner. Specifically, this phase is responsible for handling a multi-class classification task. Given a logging statement $s_{original}$ with its associated context (i.e., code snippet), the goal is to determine whether $s_{original}$ has a specific type of defect. We resolve the problem through the following steps:
@@ 3.3.2. Similarity-based multiple classification
3.4. Updating Detected Defective Logging Statement
Following the defect classification phase, where potential defective logging statements are identified, we proceed with a multi-expert collaboration framework to repair these statements, inspired by the cognitive synergy (Goertzel, 2009). In particular, we employ two LLM-based experts: a checker and an updater. In Fig. 3 and Fig. 4, we see a representation of the collaborative dynamics between two domain experts. These figures illustrate not only the specialized responsibilities assigned to each expert but also the corresponding prompts tailored for them. Our approach to designing prompts for the updating component is meticulously informed by well-established best practices in LLM prompting. This involves a thoughtful task-specific structuring and the application of few-shot learning, utilizing historical examples which have demonstrated enhancements in task performance in existing literature (Gao et al., 2024; Wei et al., 2022). Initially, the checker plays a crucial role by verifying the classification outcomes generated by the detector. Subsequently, it employs LLMs to augment these results with additional semantic insights. Following this, the role of the updater is to diligently execute updates to logging statements in accordance with the semantic guidance provided by the checker.
@@ 3.4.1. Checker
Figure 3. The Checker framework on a static-dynamic inconsistency case.
The role of checker is to reduce the false positives brought by the similarity-based detector, where non-defective logging statements could be mistakenly marked as requiring updates due to the extremely imbalanced data in real-world situations. To address this issue, we have designed a checker with two objectives. The first objective is to determine whether the detector’s decisions on updating logging statements are necessary, aiming to eliminate false positives. The second objective is to leverage the extensive knowledge of the LLM to offer additional semantic insights, thereby enhancing the effectiveness of subsequent experts.
@@ 3.4.2. Updater
We utilize log-centric changes as supplementary knowledge to guide the LLM in learning the past update patterns from maintainers. We apply a defect type-aware strategy to select proper historical LCC examples for the updater’s reference. In specific, for defect types sensitive to project characteristics (e.g., temporal relation inconsistency and readability issues), we exclusively select the LCCs from the same project to maintain the consistency in the project’s logging style (Li et al., 2024a). In contrast, for defects related to semantic information inconsistency that are unaffected by logging style and inter-project variations, we incorporate LCCs from multiple projects for the updater. We employ the BM25 (Robertson et al., 2009) ranking algorithm to identify the LCC that are the most similar to the target logging statements.
Figure 4. The Updater framework on a static-dynamic inconsistency case.
@@ 4.1. Fine-tuning Dataset
To develop a high-quality dataset for the detector fine-tuning process, our initial step involves the random selection of 20,000 cases that include logging statements, subsequently dividing them into four distinct categories. As a second step, the method detailed in Section 3.2 is utilized to produce 5,000 defective instances for each categorized type. Logging statements sourced from repositories that are meticulously maintained are assumed to be devoid of defects due to their compliance with rigorous community standards and comprehensive maintenance practices. Specifically, we select well-maintained repositories from the target repositories based on the following criteria:
@@ 4.2. Implementation Details
During LLM-based defective logging statement synthesis, we use the GPT-4o (OpenAI, 2024) API provided by Openai as the default LLM model. We use UniXcoder-base as the tokenization and embedding tool. We use Adam (Kingma and Ba, 2017) as the optimizer and set the learning rate to 5e-5, Adam epsilon to 1e-8, and dropout rate to 0.1 when fine-tuning the similarity-based classification model. We fine-tune the model for 10 epochs and set the max token number to 1024. We run them on a server with Dual Intel(R) Gold 6326 CPU, 512GB physical memory, and 4x NVIDIA RTX 3090 GPU. The OS version is Ubuntu 18.04. For all experiments related to LLM inference, we set the temperature to 0 so that LLM would generate the same output for the same query to ensure reproducibility.
@@ 5. Evaluation
We evaluate LogUpdater by answering the following research questions:
RQ1: How effective is LogUpdater at detecting and enhancing defective logging statements?
RQ2: How effective is LogUpdater at end-to-end detect and update defective logging statements?
RQ3: How do different components affect the performance of LogUpdater?
RQ4: How generalizable is LogUpdater for different LLM backbones?
RQ5: How is the practitioners’ feedback on LogUpdater?
@@ 5.1. Evaluation Dataset
To evaluate the performance of LogUpdater, we conducted experiments on three distinct datasets.
@@ 5.1.1. Existing logpatches
In our study, we employ the dataset annotated during the manual investigation presented in (2.2) as a source of existing logpatches. These consist of 258 defective instances derived from the manually categorized dataset. Furthermore, we enhance our dataset by sampling an additional 258 logging statements that are defect-free from repositories known for their meticulous maintenance as outlined in Section 4.1. This procedure results in a comprehensive dataset comprising a total of 516 cases. It is important to highlight that these specific data points are deliberately excluded from the training dataset.
@@ 5.1.3. New repositories
To evaluate the effectiveness of LogUpdater in addressing previously unidentified defects, we randomly sampled 100K log-code pairs from the target repositories but outside the collected well-maintained dataset for human evaluation. These samples were deliberately excluded from all previous processes, thereby guaranteeing that the evaluation was conducted on entirely fresh data.
@@ 5.2. Metrics
We evaluate LogUpdater from two perspectives: the detection ability and updating ability.
5.2.1. Evaluation metrics for detection ability
For the detection ability of LogUpdater, we use Precision, Recall, and F1-macro to measure multiple classification performance. Formally, $Precision=\frac{TP}{TP+FP}$, $Recall=\frac{TP}{TP+FN}$, $F_{1}=2*(\frac{Precision*Recall}{Precision+Recall})$, and $F_{1}^{macro}=\frac{1}{N}\sum_{i=1}^{N}F_{1}^{i}$. We use $F_{1}^{macro}$ to give equal weight to each class, avoiding the issue of inflating the weight of the class with more instances.
5.2.2. Evaluation metrics for updating ability
For the updating ability, we use similar metrics as logging statement generation task. Different metrics are used to evaluate different parts of the logging statements.
(1) Static content. In alignment with recent research efforts (Li et al., 2023b; Mastropaolo et al., 2022; Ding et al., 2022), the quality of generated logging texts is evaluated using two established metrics from the field of machine translation: BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004). These n-gram-based metrics quantify the similarity between the generated log messages and those crafted by developers, producing a normalized score ranging from 0 to 1, where higher scores denote greater textual correspondence. Specifically, we utilize variations of these metrics, namely BLEU-K (where K = 1, 2, 4) and ROUGE-K (where K = 1, 2, L), to measure the overlap in terms of K-grams between the generated and actual log texts.
(2) Dynamic variables. Following the recent work (Li et al., 2023b), we utilize special Precision, Recall, and F1 to assess the accuracy of updated logging variables . For each logging statement, let $S_{ud}$ represent the variables in the LogUpdater updates, and $S_{gt}$ denote the variables in the actual logging statement. We calculate and report the proportion of correctly updated variables ($precision$=$\frac{S_{ud}\cap S_{gt}}{S_{ud}}$), the proportion of actual variables that are correctly predicted by the model ($recall$=$\frac{S_{ud}\cap S_{gt}}{S_{gt}}$), and their harmonic mean ($F_{1}$=$2*\frac{Precision*Recall}{Precision+Recall}$).
@@ 5.2.2. Evaluation metrics for updating ability
Table 3. The detection result on existing logpatches and synthetic data
Existing logpatches Synthetic data
Precision Recall F1 Precision Recall F1
CodeT5+ 0.314 0.062 0.096 0.366 0.075 0.115
Claude3.5-sonnet 0.441 0.436 0.413 0.575 0.492 0.530
Deepseek-coder-v2 0.406 0.345 0.351 0.648 0.479 0.512
Deepseek-R1 0.381 0.426 0.391 0.419 0.564 0.446
GPT3.5 0.256 0.204 0.167 0.474 0.258 0.270
GPT4o 0.468 0.401 0.410 0.512 0.467 0.476
LogUpdater 0.725 0.552 0.625 0.925 0.673 0.776
5.3. RQ1: Effectiveness of LogUpdater
Table 4. The static texts updating results on existing logpatches and synthetic data
BLEU-1 BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
Existing logpatches
GPT4o-Instruct 0.690 (4.90$\uparrow$) 0.587 (12.86$\uparrow$) 0.268 (0.94$\uparrow$) 0.746 (13.01$\uparrow$) 0.593 (17.44$\uparrow$) 0.745 (12.96$\uparrow$)
GPT4o-ICL 0.681 (2.14$\uparrow$) 0.588 (13.08$\uparrow$) 0.262 (0.14$\uparrow$) 0.750 (14.38$\uparrow$) 0.603 (19.47$\uparrow$) 0.748 (13.99$\uparrow$)
GPT4o-DET 0.701 (8.28$\uparrow$) 0.620 (19.83$\uparrow$) 0.327 (8.93$\uparrow$) 0.781 (25.00$\uparrow$) 0.650 (29.01$\uparrow$) 0.779 (24.57$\uparrow$)
LogUpdater 0.799 (38.34$\uparrow$) 0.721 (41.13$\uparrow$) 0.384 (16.64$\uparrow$) 0.849 (48.28$\uparrow$) 0.738 (46.85$\uparrow$) 0.848 (48.12$\uparrow$)
Synthetic data
GPT4o-Instruct 0.546 (-7.32$\downarrow$) 0.465 (-7.86$\downarrow$) 0.316 (-2.54$\downarrow$) 0.634 (-6.08$\downarrow$) 0.499 (-9.15$\downarrow$) 0.632 (-6.35$\downarrow$)
GPT4o-ICL 0.545 (-7.56$\downarrow$) 0.468 (-7.25$\downarrow$) 0.313 (-3.00$\downarrow$) 0.639 (-4.64$\downarrow$) 0.508 (-7.19$\downarrow$) 0.637 (-4.91$\downarrow$)
GPT4o-DET 0.640 (14.89$\uparrow$) 0.579 (15.12$\uparrow$) 0.371 (5.69$\uparrow$) 0.715 (17.39$\uparrow$) 0.636 (20.69$\uparrow$) 0.714 (17.34$\uparrow$)
LogUpdater 0.659 (19.39$\uparrow$) 0.592 (17.74$\uparrow$) 0.436 (15.44$\uparrow$) 0.753 (28.40$\uparrow$) 0.648 (23.31$\uparrow$) 0.751 (28.03$\uparrow$)
Table 5. The dynamic variables updating result on existing logpatches and synthetic data
@@ 5.3. RQ1: Effectiveness of LogUpdater
GPT4o-Instruct 0.272 (0.13$\uparrow$) 0.256 (1.59$\uparrow$) 0.264 (1.15$\uparrow$)
GPT4o-ICL 0.300 (3.97$\uparrow$) 0.307 (21.82$\uparrow$) 0.304 (16.47$\uparrow$)
GPT4o-DET 0.315 (6.03$\uparrow$) 0.296 (17.46$\uparrow$) 0.305 (16.85$\uparrow$)
LogUpdater 0.439 (23.04$\uparrow$) 0.451 (26.60$\uparrow$) 0.445 (24.90$\uparrow$)
@@
GPT4o-Instruct 0.696 (14.85$\uparrow$) 0.707 (42.20$\uparrow$) 0.702 (32.57$\uparrow$)
GPT4o-ICL 0.683 (11.20$\uparrow$) 0.664 (33.73$\uparrow$) 0.673 (26.02$\uparrow$)
GPT4o-DET 0.752 (30.53$\uparrow$) 0.761 (52.85$\uparrow$) 0.756 (44.79$\uparrow$)
LogUpdater 0.795 (42.57$\uparrow$) 0.778 (56.21$\uparrow$) 0.787 (51.80$\uparrow$)
Approach. The effectiveness of LogUpdater should be evaluated from two aspects: first, the detection capability, specifically the capability to identify defective logging statements; second, the updating capability, which pertains to the similarity between the revised logging statements (i.e., prediction) and the actual ground truth (i.e., repository update history).
To evaluate detection capabilities, we employ a CodeT5+ (Wang et al., 2023) model pre-trained following the CMI-Finder (Bouzenia and Pradel, 2023) methodology as a baseline for comparison. This model has demonstrated effective performance in binary classification tasks concerning message-condition inconsistency. We adhered to the same pretraining methodology as our proposed detector, utilizing an identical fine-tuning dataset consisting of 32K samples, conducted fine-tuning over 10 epochs, and capped the maximum token length at 1024. Furthermore, given the progress in the development of LLM and their potential use in similar tasks, we also consider serval prominent LLMs as baselines, including gerneral-purpose LLMs (Claude3.5-sonnet, GPT3.5, GPT4o), code-based LLMs (Deepseek-coder-v2), and reasoning LLMs (Deepseek-R1). For the LLM-based baselines, we provide five fixed examples for task demonstration. The prompt we use for the baselines can be found in (Logupdater, 2025a).
@@
Result. Table 3 shows the detection results with different classification models. The experimental results demonstrate that LogUpdater significantly outperforms all baseline models in detecting defective logging statements across both existing logpatches and synthetic datasets. With F1 scores of 0.625 and 0.776 respectively, LogUpdater achieves substantial improvements over the next best performers Claude3.5-sonnet. This superior performance is particularly evident in precision metrics, where LogUpdater reaches 0.725 and 0.925 across the two datasets, indicating its high reliability in correctly identifying defective logging statements. General-purpose LLMs like Claude3.5-sonnet and GPT4o show moderate effectiveness (F1 scores around 0.41-0.47), while code-specialized models like Deepseek-coder-v2 perform comparably on synthetic data but lag on real-world logpatches. Furthermore, all models perform better on synthetic data than on existing logpatches, suggesting that real-world logging defects present more complex patterns than synthetically generated ones. These findings validate the effectiveness of our specialized approach for detecting logging inconsistencies in production code.
@@
Table 4 and Table 5 illustrate the updating performance on static content and dynamic variables, respectively. Compared to the three baselines, LogUpdater demonstrates significant performance improvements across various metrics on two datasets. Specifically, for GPT4o-Instruct and GPT4o-ICL, which lack defect information, the improvement on the logpatches dataset is minimal (0.14$\uparrow$ IC on BLEU-4), and there is even a regression (3.00$\downarrow$) on synthetic data. However, GPT4o-DET, which includes defect information, shows a certain level of improvement, highlighting the importance of defect information.